LASER SPECTROSCOPY OF THE X̃^{2}Π_{g}, Ã^{2}Π_{u} AND B̃^{2}Σ_{u}^{+} STATES OF BS_{2}: RENNER-TELLER, SPIN-ORBIT and K-RESONANCE EFFECTS

DENNIS J. CLOUTHIER and SHENG-GUI HE, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; ALLAN G. ADAM and \underline{DENNIS W. TOKARYK}, Departments of Chemistry and Physics, and the Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, Canada, E3B 6E2.

The lowest-lying vibronic levels of the X̃, Ã, and B̃ states of BS_{2} have been investigated at high resolution using a combination of room-temperature absorption and supersonic jet data. In both cases, the BS_{2} radical was prepared in an electric discharge using a precursor gas mixture of BCl_{3}, CS_{2}, and either helium or argon. Extensive absorption spectra were obtained for the 0_{0}^{0} and 2_{1}^{1} bands of the Ã^{2}Π_{u} − X̃^{2}Π_{g} electronic transition in the visible using a mutipass discharge cell and a scanning ring dye laser. The Ã − X̃ and B̃ − X̃ 2_{1}^{1} bands of jet-cooled BS_{2} were also studied with laser-induced fluorescence techniques. By fitting the 0_{0}^{0} bands of both electronic transitions simultaneously, we were able to precisely determine the spin-orbit splittings in both the Ã and X̃ states. The 2_{1}^{1} bands were fitted in a merged analysis in order to determine the relative separations of the vibronic components of the ground and first excited state bending levels as accurately as possible. Due to a large spin-orbit splitting and small Renner-Teller interaction, the Ã state bending level shows small but definite K-resonance effects, which were fitted using a full matrix for the four components of v_{2} = 1. The resulting parameters were used along with previously published data to refine the Renner-Teller analyses in both the Ã^{2}Π_{u} and X̃^{2}Π_{g} electronic states.